Dad is 37 years older than his son Joel.Dad is 4 years younger than Mum.The total of their ages added together is 99.How old is Mum?

Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52781)

.

Mum + Dad   + Joel       = 99,   or


  x + (x-4) + ((x-4)-37) = 99.


Simplify and solve


  3x - 45 = 99

  3x      = 99 + 45

   x      = 33 + 15 = 48.     Mum's age.     ANSWER


Answer by josgarithmetic(39617)

Dad           m-4
Joel          (m-4)-37
Mum            m
TOTAL          99